OpenStudy (anonymous):

Revision: I can solve these questions but occasionally get stuck, can someone point out the easiest method to check for one-one, onto? f:[-1,1] -> R given by f(x) = x/x + 2 is one-one Find the inverse of the function f: [-1,1] -> range f @rational

OpenStudy (rational):

y = x/(x+2) step1 : swap x and y x = y/(y+2) step2 : solve y x(y+2) = y xy+2x = y 2x = y(1-x) 2x/(1-x) = y so the inverse is 2x/(1-x)